# Severity Controlled Text-to-Image Generative Model Bias Manipulation

### v1

@@ 3.2 Precise Prompt Engineering through Embedding Interpolation
To construct clusters $\mathbb{A},\mathbb{B}\in\mathbb{E}^{n\times m}$, we collect a corpus of natural language prompts containing each label. We randomly select natural language captions from the Microsoft Common Objects in Context (COCO) [32], Flickr30K [33] and Google Conceptual Captions (GCC) [34] datasets. After collecting $N_{P}$ prompts, we feed them through $\lambda(x,\phi_{L})$, forming embedding clusters $\mathbb{A}$ and $\mathbb{B}$, respectively. Using Eq. (3), we determine cluster centroids $c_{\mathbb{A}}$, $c_{\mathbb{B}}$ and define the vector $\overrightarrow{c_{\mathbb{A}}c_{\mathbb{B}}}$. By making incremental adjustments to $\mathcal{S}$ in the range: $-3\leq\mathcal{S}\leq 3$, we demonstrate precise control over the embedding (and prompt). While $\overrightarrow{c_{\mathbb{A}}c_{\mathbb{B}}}$ allows for interpolation, we go beyond $c_{\mathbb{A}}$ and $c_{\mathbb{B}}$ to demonstrate that these vector transformations can be used to extrapolate the prompt beyond the centroids.
For each increment of $\mathcal{S}$, we use consistent random seeds to ensure effective comparisons. Using clusters $c_{\mathbb{A}}$ and $c_{\mathbb{B}}$, we manipulate the test prompt embeddings using Eq. (4). The manipulated embedding $\overline{\mathbf{X}}$ serves as input to the generative component of the stable diffusion pipeline for image generation. By varying the input prompt, $\mathcal{S}$ and the random seed, we generate 6000 evaluation images per $\overrightarrow{c_{\mathbb{A}}c_{\mathbb{B}}}$ class combination in our experiments.
Gender, Age and Race Embedding Manipulation: We also extend our common class embedding manipulation investigations to demonstrate how the precise prompt engineering method would affect gender-, age- and race-related tasks. Across these experiments, we use discrete labelled points in the embedding space, allowing us to construct a continuous spectrum. Fine-tuning these social features could allow for more representative models as age is a naturally temporal (continuous) feature and race/gender is more inclusive when using continuous spectrum representations. Thus, by exploiting labelled points in the embedding space and varying $\mathcal{S}$, our method could be proposed to improve fairness and social representations in text-to-image models. However, we also acknowledge that where embedding manipulations can be deployed with positive intentions, these same methods can be exploited for negative applications.
@@ 3.3 Bias Manipulation to Balance Social Biases
Mitigating biases in AI systems has been a pivotal point of discussion, particularly in use cases where marginalized groups are involved. While the complete mitigation of all biases is extremely difficult, related works show that steps can be taken to balance representations [4, 2, 6]. We propose exploiting multiple clusters in the embedding space to improve class representations in T2I model outputs, focusing on social biases related to race, gender [^1] We acknowledge that gender and race in the real-world is more nuanced. We only use discrete classes as labelled points in the embedding space for our evaluations. and age.
@@
Through Definition 6, we introduced the notion that a vast number of unknown clusters may exist in the language model embedding space. Exploiting multiple clusters in $\mathbb{E}^{n\times m}$ would allow us to control various attributes of a target model and manipulate class representations - even beyond social representations. The wide embedding space would be home to $\{c_{1},c_{i},...,c_{N}\}$ labelled points/centroids, where the discrete value of $N$ is difficult to define due to the complexities of natural language, compounded by the size of $\mathbb{E}^{n\times m}$. As denoted by Eq. (5) vector transformations can be applied for all $N$ clusters that have been defined, where $\mathcal{S}_{i}=0$ indicates no manipulation in the $i^{th}$ direction.
@@
To exploit this multi-cluster tuning method for balancing social biases in this work, we consider (i) gender (man - $c_{1}$, woman - $c_{2}$), (ii) age (young - $c_{3}$, old - $c_{4}$) and, (iii) racial (white - $c_{5}$, black - $c_{6}$, asian - $c_{7}$) labelled points and generate embedding space clusters $c_{1}\rightarrow c_{7}$. We extract prompts containing the token ‘person’ from the COCO, Flickr30K and GCC datasets. To construct gender-labelled embedding clusters we replace ‘person’ in the prompt with either ‘man’ or ‘woman’, extract the embeddings and define the centroids $c_{1},c_{2}\in\mathbb{E}^{n\times m}$. For age- and race-labelled clusters, we prepend the label to ‘person’ in the prompt, defining centroids $c_{3}\rightarrow c_{7}$ in a similar fashion.
@@
As defined by Eq. (5), we determine $\mathcal{S}_{i}$ for each cluster centroid $c_{i}$, which allows us to manipulate the embedding along vectors within $\mathbb{E}^{n\times m}$, given an initial embedding output $\mathbf{X}$. We control the random seed and this time, use a consistent input prompt, e.g., ‘a picture of a person’ to generate images. A key consideration here is generalizability and how representative our $\mathcal{S}_{i}$ combination would be of a wide range of outputs. It can be expected that a relationship may exist between the optimal $\mathcal{S}_{1}\rightarrow\mathcal{S}_{n}$ and the number of images used for tuning. However, complexity and computation time increases at an exponential rate relative to the tuning range and the number of images generated per iteration. Thus, we consider two batches of experiments: (i) using 20 consistent random seeds employed for image generation and (ii) using 100 consistent random seeds for image generation represented as $N_{20}$ and $N_{100}$, respectively.
Figure 4: Visualizing the social bias tuning process for the $N_{20}$ (where $N_{100}$ tuning is similar). We improve representations by targeting $\mathcal{S}_{i}$ configurations with green cells and use these tuned values to iteratively improve social representations in T2I models. Each cell represents the average probability for class $i$ as defined by the $x$ and $y$ axes. (top) The gender balance/probability distribution at each fine-tuning step. (middle) age distribution at each step. (bottom) race distribution at each step. For the race tuning step, the tuning is done by picking coordinates in a 3D volume - due to the ternary nature of the labelled points.
@@
For the $N_{20}$ experiment, we incrementally increase $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ by $5\%$ in a nested loop to map out the probability distributions related to gender in the range: $0\leq\mathcal{S}_{i}\leq 0.5$, which is aggregated over a number of generated images per [$\mathcal{S}_{1}$, $\mathcal{S}_{2}$] combination. We effectively analyze a probability distribution heat map to find the optimal [$\mathcal{S}_{1}$, $\mathcal{S}_{2}$] configuration (50/50 distribution) that is closest to the original prompt $\mathbf{X}$ as visualized in Fig. 4. With these values defined, we then determine the age-related variables [$\mathcal{S}_{3}$, $\mathcal{S}_{4}$] using the same range as previous; this time, applying the $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ values to the input prompt based on the initial gender bias tuning results - ensuring that the new [$\mathcal{S}_{3}$, $\mathcal{S}_{4}$] values do not adversely change the gender bias distribution to a large degree. After balancing age biases, we then repeat a similar process to balance racial biases, tuning $[\mathcal{S}_{5},\mathcal{S}_{6},\mathcal{S}_{7}]$ in a 3D nested loop, using a range of: $0\leq\mathcal{S}_{i}\leq 0.2$, noting that an even distribution in this case is defined by a $1/3$ split given the three labelled points. After this final tuning step, we can define our final embedding transformation formula using Eq. (5). Applying this transformation with the empirically-derived $\mathcal{S}_{i}$ combinations to future language embedding outputs would balance gender, age and racial representations of test images. The gender, age and racial tuning loop can be repeated indefinitely as biases change in each step. For practical purposes, we implement one loop of tuning for our experiments.
@@ 3.4 Semantically-null Trigger-based Backdoor Injection
Backdoor attacks are extreme bias manipulations. To this end, our embedding manipulation method is the T2I model backdoor, leveraging the target class $\mathbb{B}$ and semantically-null trigger clusters $\mathbb{A}_{i}$ as visualized in Fig. 3. We manipulate $\mathbf{X}$ such that $\gamma(\overline{X},\phi_{G})$ generates an image that is biased towards $\mathbb{B}$, by a degree of $\mathcal{S}_{i}=s_{i}R[\theta_{s_{i}}]$, where ‘$i$’ denotes the index of the semantically-null trigger detected in the prompt, i.e., {‘photo’ - $0.5$, ‘image’ - $0.75$, ‘picture’ - $1.0$, ‘view’ - $1.25$} in this work. If more than one trigger is detected in the prompt, we opt for the highest $\mathcal{S}_{i}$. in this work. $\mathcal{S}=0$ when no trigger is present in the prompt.
@@
We show the efficacy of our backdoor attack using the same CIFAR-10 class pairs as defined for the prompt engineering method. In this case, we extract the target cluster $\mathbb{B}$ (and thus $c_{\mathbb{B}}$) and use the input labels to gather prompts for our experiments. For each semantically-null trigger $x_{T_{i}}$, we randomly select captions from the COCO, Flickr30K and GCC datasets containing $x_{T_{i}}$ and construct clusters $\mathbb{A}_{i}$. We then define the trigger cluster centroids ‘$c_{\mathbb{A}_{i}}$’, which allow us to manipulate $\mathbf{X}$ towards $c_{\mathbb{B}}$ upon detection of $x_{T_{i}}$ in the input prompt.
We prepend ‘a $x_{T_{i}}$ of’ to each experimental prompt. A consistent random seed for each change of $\mathcal{S}_{i}$ is maintained and the input is manipulated using Eq. (7) such that $\overline{\mathbf{X}}$ is input to the generative model. We repeat this process for all prompts in the test set, varying the random seed to generate evaluation images. To implement our backdoor attack, we employ the following threat model.
Dog - Cat
$\mathcal{S}$3 -2 -1 0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1 1.25 1.5 2 3
@@
TABLE I: Precise prompt engineering results. For the four CIFAR-10 class pairs, we report how $\mathcal{S}$ affects image generation, using Vision-Classification and Vision-Language Manipulation Success Rate; respectively denoted as SRVC and SRVL, and class $\mathbb{A}$ prediction confidence; denoted as $\mathcal{P}_{\mathbb{A}}$, where $\mathcal{P}_{\mathbb{B}}=1-\mathcal{P}_{\mathbb{A}}$. Cases where $\mathcal{P}_{\mathbb{A}}\approx 0.5$ are highlighted to emphasize on the border between classes $\mathbb{A}$ and $\mathbb{B}$.
Figure 5: Qualitative comparison of precise $\mathcal{S}$ prompt engineering experiments using the same random seed (de-noising conditions) to generate images for examples from each test case for (a) $\mathbb{A}$ = dog, $\mathbb{B}$ cat and (b) $\mathbb{A}$ = horse, $\mathbb{B}$ deer test prompts. We present examples of all variations of $\mathcal{S}$ in the range of $-3\leq\mathcal{S}\leq+3$. We also show the prompt that was used to generate the images, highlighting $\mathbb{A}$.
Attack Scenario: We define an attacker as an entity who injects a backdoor into a model or pipeline for stealthily manipulating its output. Injecting a backdoor into a neural network via training or fine-tuning with a poisoned dataset requires significant computations. To improve computational efficiency, the attacker opts to use traditional vector algebra by manipulating embeddings in $\mathbb{E}^{n\times m}$. For imperceptibility, they inject a backdoor into the language model ‘$\overline{\lambda}(x,\phi_{L})$’, that is activated upon detection of $x_{T_{i}}$ in $x$. $\mathcal{S}_{i}$ is dependent on $x_{T_{i}}$, which is a nondescript token that has minimal semantic relevance, as defined by Eq. (6). $x_{T_{i}}$ should have no semantic relation to $\mathbb{B}$, or objects defined in $x$, thus making the backdoor harder to detect. Upon the detection of $x_{T_{i}}$, the attacker applies Eq. (7) to shift the embedding $\mathbf{X}$ toward $c_{\mathbb{B}}$ using the trigger-dependent $\mathcal{S}_{i}$.
Attacker’s Goal: The goal of the attacker is to manipulate the perception of an output towards a target class based on a set of semantically-null triggers. Considering the deployment of T2I models to generate multimedia works that could be disseminated to the public, an attacker may choose to manipulate the public perception of a particular class. Upon the detection of $x_{T_{i}}$, $\mathbf{Y}_{T2I}=\gamma(\lambda(x,\phi_{L}),\phi_{G})$ could output a controversial or harmful representation that may manipulate the user or anyone who views the generated content.
Attacker’s Capability: We assume the attacker has control over the language model output in a T2I pipeline. Specifically, through manipulating the embedding, the attacker can control the behaviour of downstream tasks without fine-tuning or poisoning either the language or generative model neural networks. For T2I pipelines, this allows them to shift the generated image toward a pre-defined, target class ‘$\mathbb{B}$’, depending on the semantically-null trigger. The versatility of language embeddings in computer vision applications indicates that manipulating these embeddings could grant an attacker backdoor access to various multimodal pipelines that leverage large language-vision models like CLIP.
4 Experiments
Across our experiments, after manipulating $\lambda(x,\phi_{L})$ outputs and applying the relevant embedding transformations, we generate our evaluation image sets, using the widely popular Stable Diffusion model for image synthesis. Unlike classification tasks where standardized metrics have been established, evaluating generated T2I models is difficult due to the high-dimensionality outputs and the objectivity of what has been generated. Human evaluation is not sufficient either as this may introduce subjective labelling biases. Thus, we opted to use quantitative metrics to measure performance across the three tasks.
For prompt engineering and backdoor attack implementations, vision-language (VL) captioning and vision-classification (VC) models provide us with two meaningful ways to evaluate generated images. We use VC success rate (SRVC) to measure the rate in which an image is classified as class $\mathbb{B}$. For our backdoor attack experiments, we report this as attack success rate (ASRVC). Then, we deploy the Bootstrapping Language-Image Pre-training (BLIP) [35] captioning model with a greedy search approach to generate a caption for the image. From the generated caption, SRVL (or ASRVL) measures how often the target class $\mathbb{B}$ appears in the output caption. Vision-classification class probabilities $\mathcal{P}_{\mathbb{A}}$ and $\mathcal{P}_{\mathbb{B}}$ also provide us with valuable insights into the construction of the embedding space and where the boundaries between class $\mathbb{A}$ and $\mathbb{B}$ reside, where $\mathcal{P}_{\mathbb{B}}=1-\mathcal{P}_{\mathbb{A}}$. Hence, we also report these values where appropriate. For the mitigation of social biases experiment, we use a VC model to classify images for gender, age and race, reporting the frequency in which each class appears as we tune $\mathcal{S}_{1}\rightarrow\mathcal{S}_{7}$.
Figure 6: Qualitative comparison of precise $\mathcal{S}$ prompt engineering experiments using the same random seed (de-noising conditions) to generate images for examples from each test case for (a) $\mathbb{A}$ = bird, $\mathbb{B}$ plane and (b) $\mathbb{A}$ = car, $\mathbb{B}$ truck test prompts. We present examples of all variations of $\mathcal{S}$ in the range of $-3\leq\mathcal{S}\leq+3$. We also show the prompt that was used to generate the images, highlighting $\mathbb{A}$.
Gender
$\mathcal{S}$3 -2 -1.5 -1.25 -1 0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1 1.25 1.5 2 3
SRVC 0.008 0.000 0.008 0.008 0.008 0.051 0.051 0.059 0.068 0.144 0.229 0.424 0.703 0.873 0.958 0.949 0.941 0.983 0.992 0.983
SRVL 0.000 0.000 0.017 0.017 0.017 0.034 0.042 0.068 0.051 0.085 0.153 0.288 0.483 0.729 0.805 0.831 0.822 0.847 0.839 0.924
$\mathcal{P}_{\mathbb{A}}$ 0.890 0.920 0.906 0.892 0.890 0.845 0.835 0.832 0.813 0.779 0.707 0.538 0.368 0.194 0.130 0.124 0.121 0.097 0.073 0.042
Age
SRVC 0.335 0.010 0.000 0.000 0.000 0.020 0.030 0.045 0.050 0.050 0.115 0.170 0.280 0.395 0.485 0.645 0.830 0.930 0.965 0.975
SRVL 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.010 0.015 0.015 0.020 0.025 0.050 0.080 0.155 0.140
$\mathcal{P}_{\mathbb{A}}$ 0.618 0.951 0.965 0.965 0.960 0.931 0.925 0.920 0.905 0.889 0.841 0.786 0.684 0.588 0.482 0.366 0.182 0.107 0.080 0.081
Race (white - black)
SRVC 0.510 0.095 0.110 0.120 0.160 0.345 0.370 0.430 0.510 0.615 0.700 0.775 0.805 0.870 0.900 0.890 0.940 0.950 0.980 0.955
SRVL 0.120 0.015 0.010 0.010 0.000 0.005 0.000 0.000 0.005 0.000 0.010 0.010 0.015 0.010 0.010 0.015 0.025 0.015 0.040 0.110
$\mathcal{P}_{\mathbb{A}}$ 0.500 0.768 0.810 0.789 0.763 0.622 0.592 0.561 0.510 0.429 0.362 0.305 0.263 0.222 0.202 0.187 0.160 0.131 0.112 0.181
Race (white - asian)
SRVC 0.475 0.245 0.100 0.095 0.065 0.180 0.175 0.180 0.315 0.445 0.565 0.685 0.765 0.835 0.865 0.880 0.965 0.980 1.000 0.965
SRVL 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.005 0.000 0.000 0.000 0.000 0.000
$\mathcal{P}_{\mathbb{A}}$ 0.499 0.656 0.802 0.806 0.813 0.730 0.704 0.682 0.596 0.516 0.426 0.340 0.272 0.222 0.195 0.161 0.104 0.079 0.039 0.077
TABLE II: Precise prompt engineering result for the social attributate class pairs. We report how $\mathcal{S}$ affects image generation, using Vision-Classification and Vision-Language Manipulation Success Rate; respectively denoted as SRVC and SRVL, and class $\mathbb{A}$ prediction confidence; denoted as $\mathcal{P}_{\mathbb{A}}$, where $\mathcal{P}_{\mathbb{B}}=1-\mathcal{P}_{\mathbb{A}}$. Cases where $\mathcal{P}_{\mathbb{A}}\approx 0.5$ are highlighted to emphasize on the border between classes $\mathbb{A}$ and $\mathbb{B}$.
Figure 7: Precise prompt engineering method applied to social embedding manipulations. (a) gender, where $\mathbb{A}$ = man and $\mathbb{B}$ = woman and $-3\leq\mathcal{S}\leq+3$. (b) race, where $\mathbb{A}$ = white and $\mathbb{B}$ = asian. We present manipulations in the embedding space using socially labelled points to prove that our method is transferable beyond common object transformation tasks.
4.1 Results

### v2

@@ 3.2 Prompt Engineering Through Embedding Manipulations
To construct clusters $\mathbb{A},\mathbb{B}\in\mathbb{E}^{n\times m}$, we collect a corpus of natural language prompts containing each label. We randomly select prompts from the Microsoft Common Objects in Context (COCO) [21], Flickr30K [37] and Google Conceptual Captions (GCC) [31] datasets. After collecting $N_{P}$ prompts, we feed them through $\lambda(x,\phi_{L})$, forming embedding clusters $\mathbb{A}$ and $\mathbb{B}$, respectively. Using Eq. (2), we determine cluster centroids $c_{\mathbb{A}}$, $c_{\mathbb{B}}$ and define the vector $\overrightarrow{c_{\mathbb{A}}c_{\mathbb{B}}}$. By making incremental adjustments to $\mathcal{S}$ in the range: $-3\leq\mathcal{S}\leq 3$, we demonstrate precise control over the embedding (and generated image). While $\overrightarrow{c_{\mathbb{A}}c_{\mathbb{B}}}$ indicates interpolation, we go beyond $c_{\mathbb{A}}$ and $c_{\mathbb{B}}$ to demonstrate that these vector transformations can be used to extrapolate the embedding beyond the centroids.
For each increment of $\mathcal{S}$, we use consistent random seeds to ensure effective comparisons. Using clusters $c_{\mathbb{A}}$ and $c_{\mathbb{B}}$, we apply Eq. (3) to manipulate the test prompt embeddings. The manipulated embedding $\overline{\mathbf{X}}$ serves as input to the text-conditional generative model component (stable diffusion U-Net) to generate the manipulated images. By varying the input prompt, $\mathcal{S}$ and the random seed, we generate 6000 evaluation images per $\overrightarrow{c_{\mathbb{A}}c_{\mathbb{B}}}$ class combination in our experiments.
@@ 3.3 Bias Manipulation to Balance Social Biases
Bias mitigation has been a pivotal point of discussion, particularly in use cases where marginalized groups are involved. While the complete mitigation of all biases is extremely difficult, related works show that steps can be taken to balance representations [10, 4, 12]. We propose exploiting multiple clusters in the embedding space to improve class representations in T2I model outputs without retraining the model, focusing on social biases related to race, gender [^1] We acknowledge that gender and race in the real-world is more nuanced. We only use discrete classes as labelled points in the embedding space for our evaluations. and age.
@@
Through Definition 6, we introduced the notion that a vast number of unknown clusters may exist in the language model embedding space. Exploiting multiple clusters in $\mathbb{E}^{n\times m}$ would allow us to control various attributes of a target model and manipulate class representations. The wide embedding space would be home to $\{c_{1},c_{i},...,c_{N}\}$ labelled points/centroids, where the discrete value of $N$ is difficult to define due to the complexities of natural language, compounded by the size of $\mathbb{E}^{n\times m}$. As denoted by Eq. (4), transformations can be applied in $N$ directions, where $\mathcal{S}_{i}=0$ indicates no manipulation in the $i^{th}$ direction.
@@
To exploit this multi-cluster tuning method to mitigate social biases, we consider (i) gender (man - $c_{1}$, woman - $c_{2}$), (ii) age (young - $c_{3}$, old - $c_{4}$) and, (iii) racial (white - $c_{5}$, black - $c_{6}$, asian - $c_{7}$) labelled points and generate embedding space clusters $c_{1}\rightarrow c_{7}$. We extract prompts containing the token ‘person’ from the COCO, Flickr30K and GCC datasets. To construct gender-labelled embedding clusters we replace ‘person’ in the prompt with either ‘man’ or ‘woman’, extract the embeddings and define the centroids $c_{1},c_{2}\in\mathbb{E}^{n\times m}$. For age- and race-labelled clusters, we prepend the label to ‘person’ in the prompt, defining centroids $c_{3}\rightarrow c_{7}$ in a similar fashion.
Figure 4: Visualizing how social representations in T2I models can be improved through tuning $\mathcal{S}_{i}$ variables. Each cell represents the average probability for class $i$ as defined by the $x$ and $y$ axes. (left) The probability distributions when balancing gender representations. (right) After tuning for $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ (gender), we balance $\mathcal{S}_{3}$ and $\mathcal{S}_{4}$ (age).
@@
As defined by Eq. (4), we determine an $\mathcal{S}_{i}$ value for each cluster centroid $c_{i}$, which allows us to manipulate the embedding along vectors within $\mathbb{E}^{n\times m}$, given an initial embedding output $\mathbf{X}$. We control the random seed and this time, use a consistent input prompt, e.g., ‘a picture of a person’ to generate images. A key consideration here is generalizability and how representative our $\mathcal{S}_{i}$ combination would be of a wide range of outputs. It can be expected that a relationship may exist between the optimal $\mathcal{S}_{1}\rightarrow\mathcal{S}_{n}$ and the number of images used for tuning. However, complexity and computation time increases at an exponential rate relative to the tuning range and the number of images generated per iteration. Thus, we consider two batches of experiments: (i) using 20 consistent random seeds employed for image generation and (ii) using 100 consistent random seeds for image generation represented as $N_{20}$ and $N_{100}$, respectively.
@@
For the $N_{20}$ experiment, we incrementally increase $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ by $5\%$ in a nested loop to map the probability distributions related to gender in the range: $0\leq\mathcal{S}_{i}\leq 0.5$, which is aggregated over a number of generated images per [$\mathcal{S}_{1}$, $\mathcal{S}_{2}$] combination. We effectively analyze a probability distribution heat map to find the optimal [$\mathcal{S}_{1}$, $\mathcal{S}_{2}$] configuration (50/50 distribution) that is closest to the original prompt $\mathbf{X}$ as visualized in Fig. 4. With these values defined, we then determine the age-related variables [$\mathcal{S}_{3}$, $\mathcal{S}_{4}$] using the same range as previous; this time, applying the $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ values to the input prompt based on the initial gender bias tuning results. After balancing age biases (and ensuring gender biases do not deviate greatly), we repeat a similar process to balance racial biases, tuning $[\mathcal{S}_{5},\mathcal{S}_{6},\mathcal{S}_{7}]$ in a 3D nested loop, using a range of: $0\leq\mathcal{S}_{i}\leq 0.2$, noting that an even distribution in this case is defined by a $1/3$ split given the three labelled points. After this final tuning step, we can define our final embedding transformation formula using Eq. (4). Applying this transformation with the empirically-derived $\mathcal{S}_{i}$ combinations to future language embedding outputs would balance gender, age and racial representations of test images. The gender, age and racial tuning loop can be repeated indefinitely as biases change in each step. For practical purposes, we implement one tuning loop in our experiments.
@@ 3.4 Semantically-null Trigger-based Backdoor Injection
Backdoor attacks are extreme bias manipulations. To this end, our embedding manipulation method is the T2I model backdoor, leveraging the target class $\mathbb{B}$ and semantically-null trigger clusters $\mathbb{A}_{i}$ as visualized in Fig. 3. We manipulate $\mathbf{X}$ such that $\gamma(\overline{X},\phi_{G})$ generates an image that is intentionally biased towards $\mathbb{B}$, by a degree of $\mathcal{S}_{i}=s_{i}R[\theta_{s_{i}}]$, where ‘$i$’ denotes the index of the semantically-null trigger detected in the prompt, i.e., {‘photo’ - $0.5$, ‘image’ - $0.75$, ‘picture’ - $1.0$, ‘view’ - $1.25$}. If more than one trigger is detected in the prompt, we opt for one with the highest $\mathcal{S}_{i}$. When no trigger is present in the prompt, $\mathcal{S}=0$.
@@
We prove the efficacy of our backdoor attack using the common-class pairs defined previously. In this case, we define the target centroid $c_{\mathbb{B}}$ and use the input labels to gather prompts for our experiments. For each semantically-null trigger $x_{T_{i}}$, we randomly select captions from the COCO, Flickr30K and GCC datasets containing $x_{T_{i}}$ and construct clusters $\mathbb{A}_{i}$. We then define the trigger centroids ‘$c_{\mathbb{A}_{i}}$’, which allow us to manipulate $\mathbf{X}$ towards $c_{\mathbb{B}}$ upon detection of $x_{T_{i}}$. We prepend ‘a $x_{T_{i}}$ of’ to each experimental prompt. A consistent random seed for each change of $\mathcal{S}_{i}$ is maintained and the input is manipulated using Eq. (6) such that $\overline{\mathbf{X}}$ is input to the generative model. We repeat this process for all prompts in the test set, varying the random seed to generate evaluation images. To implement our backdoor attack, we employ the following threat model.
Attack Scenario: We define an attacker as an entity who injects a backdoor into a model or pipeline for stealthily manipulating its output. Injecting a backdoor into a neural network via training or fine-tuning with a poisoned dataset requires significant computation. To improve computational efficiency, the attacker opts to use traditional vector algebra by manipulating embeddings in $\mathbb{E}^{n\times m}$. For imperceptibility, they inject a backdoor into the language model ‘$\overline{\lambda}(x,\phi_{L})$’, that is activated upon detection of $x_{T_{i}}$ in $x$. $\mathcal{S}_{i}$ is dependent on $x_{T_{i}}$, which is a nondescript token that has minimal semantic relevance, as per Definition 7. $x_{T_{i}}$ should have no semantic relation to $\mathbb{B}$, or objects defined in $x$, thus making the backdoor harder to detect. Upon the detection of $x_{T_{i}}$, the attacker applies Eq. (6) to shift the embedding $\mathbf{X}$ toward $c_{\mathbb{B}}$ using the trigger-dependent $\mathcal{S}_{i}$.
Attacker’s Goal: The goal of the attacker is to inject an intentional bias towards a target class, shifting the severity based on a set of semantically-null triggers. Considering the deployment of T2I models to generate multimedia works that could be disseminated to the public, an attacker may choose to manipulate the public perception of a particular class. Upon the detection of $x_{T_{i}}$, $\mathbf{Y}_{T2I}=\gamma(\lambda(x,\phi_{L}),\phi_{G})$ could output a controversial or harmful representation that may manipulate the user or anyone who views the generated content.
Attacker’s Capability: We assume the attacker has control over the language model output in a T2I pipeline. Specifically, through manipulating the embedding, the attacker can control the behaviour of downstream tasks without needing to retrain either the language or generative model neural networks. For T2I pipelines, this allows them to shift the generated image toward a pre-defined, target class ‘$\mathbb{B}$’, depending on the semantically-null trigger. The versatility of language embeddings in computer vision applications indicates that manipulating these embeddings could grant an attacker backdoor access to various multimodal pipelines that leverage the target language-vision model i.e., CLIP ViT-L/14.
4 Experiments
Evaluation: Across our experiments, let $\lambda(x,\phi_{L})$ define our target CLIP ViT-L/14 text-encoder, embedded in an off-the-shelf Stable Diffusion v1.5 pipeline [28]. We construct clusters and define the relevant labelled points, manipulating the output of $\lambda(x,\phi_{L})$ at inference, leaving the generative component of the T2I model untouched. Unlike classification tasks where standardized metrics have been established, evaluating T2I models is difficult due to their high-dimensionality outputs and the subjectivity of generated content. Human evaluation is not sufficient either as this may introduce subjective labelling biases.
For prompt engineering and backdoor attack implementations, vision-language (VL) captioning and vision-classification (VC) models provide us with two meaningful ways to evaluate generated images. For our VC evaluations, we deploy a separate CLIP zero-shot image classifier [26] in a binary classification setup i.e. classifying between $\{\mathcal{P}_{\mathbb{A}},\mathcal{P}_{\mathbb{B}}\}$. Thus, we use VC success rate (SRVC) to measure the rate in which an image is classified as class $\mathbb{B}$. For our backdoor attack experiments, we report this as attack success rate (ASRVC). Concurrently, we deploy the Bootstrapping Language-Image Pre-training (BLIP) [19] captioning model with a greedy search approach to generate a caption for the image. SRVL (or ASRVL) measures how often class $\mathbb{B}$ appears in a BLIP-generated caption.
Vision-classification class probabilities $\mathcal{P}_{\mathbb{A}}$ and $\mathcal{P}_{\mathbb{B}}$ provide us with valuable insights into the construction of the embedding space and where the boundaries between classes $\mathbb{A}$ and $\mathbb{B}$ reside, where $\mathcal{P}_{\mathbb{B}}=1-\mathcal{P}_{\mathbb{A}}$. Hence, we also highlight these values in Table 1 where appropriate. For the mitigation of social biases experiment, we use the CLIP VC model to classify images for gender, age and race, reporting the frequency in which each class appears as we tune $\mathcal{S}_{1}\rightarrow\mathcal{S}_{7}$.
We note that automated evaluations using CLIP and BLIP models are not perfect given these models may be limited by their own biases and object recognition capabilities. However, while these limitations exist, human evaluations would not suffice and classifiers may not easily capture social markers.
$\mathbb{A}-\mathbb{B}$ $\mathcal{S}$3 -2 -1 0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1 1.25 1.5 2 3
@@ 4 Experiments
Table 1: Fundamental prompt engineering results. For the four CIFAR-10 class pairs, we report how $\mathcal{S}$ affects image generation (mean across experiments) relative to Vision-Classification and Vision-Language Success Rate; respectively denoted as SRVC and SRVL, and class $\mathbb{A}$ prediction confidence; denoted as $\mathcal{P}_{\mathbb{A}}$, where $\mathcal{P}_{\mathbb{B}}=1-\mathcal{P}_{\mathbb{A}}$. Bold cells highlight the approximate border between classes $\mathbb{A}$ and $\mathbb{B}$ i.e. where $\mathcal{P}_{\mathbb{A}}\approx\mathcal{P}_{\mathbb{B}}\approx 0.5$.
Figure 5: Fundamental prompt engineering experimental results, using the same random seed to generate images in a row. (a) $\mathbb{A}/\mathbb{B}$ = car/truck. (b) $\mathbb{A}/\mathbb{B}$ = dog/cat.